OpenStudy (anonymous):

The number of full time nurses employed at a local hospital increased from 550 in 2003 to 725 in 2008. Predict the number of full time nurses in 2010. please show work

OpenStudy (dumbcow):

think of years as x_values and num nurses as y_values slope = (725-550)/(2008-2003) = 175/5 = 35 so for every additional year for num nurses increases by 35 2009 is 760 2010 is 795